A sphere of mass m and radius r , Rolls and encounters an obstacle of height r/2 . What is the angular velocity of the sphere ?

Grade:

1 Answers

Vikas TU
4 years ago
Dear student 
During collision, large impulsive forces are generated at contact point A, neglecting weight of sphere.
By the conservation of angular momentum about A.
mv(r-h) + 2/5 mr^2(v/r) = mv'r + 2/5 (v'/r) 
3/4 mvr + 2/5 mvr = 7/5 mv'r 
v' = 23/28 v 
w' = v'/r = 23/28r